# Orthogonal Block Kaczmarz Algorithm Based on Preprocessing Technology

### v3

@@ Abstract
In this paper, a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques is proposed to solve large-scale sparse linear systems $Ax=f$. Firstly, the sparse Reverse Cuthill-McKee Algorithm (RCM) is used to preprocess the linear system, and then a new partitioning strategy is proposed to divide the mutually orthogonal blocks into one category, which greatly reduces computational complexity. In addition, we provide a theoretical proof of the convergence of the POBK algorithm and analyze its faster convergence. Finally, numerical simulations demonstrate that our method outperforms the latest Kaczmarz methods such as GRBK, RBK(k), GREBK(k), and aRBK in terms of CPU time.
@@
Kaczmarz algorithm, Sparse Reverse Cuthill-McKee ordering, Preprocessing techniques, Orthogonal partitioning
@@ 1 Introduction
This paper mainly considers the solution of a large sparse linear system in the form of
$Ax=f$ (1.1)
where $A\in{{\mathbb{R}}^{m\times m}}$, $f\in{{\mathbb{R}}^{m}}$, and $x\in{\mathbb{R}^{m}}$. Nowadays, one of the classic and effective iterative projection methods for solving such problems is the Kaczmarz method [1, 2].
@@
The convergence of the classic Kaczmarz method depends on the order of matrix row selection, and the convergence speed is slow and it is difficult to provide a convergence theory. Therefore, Strohmer and Vershynin [3, 4] proposed a random Kaczmarz (RK) algorithm with an expected exponential convergence rate. However, the probability criterion for randomly selecting rows in the RK algorithm is not optimal. Based on the idea of greed, Ansorge proposed the greedy residual Kaczmarz (GEK) method [5]. Nutini et al. proposed the greedy distance Kaczmarz (GDK) method [6]. Bai Zhongzhi and Wu Wenting combined the mathematical concept of randomness and introduced a new and more effective selection criterion by defining an indicator set. They proposed the greedy random Kaczmarz (GRK) algorithm and its relaxed version [7, 8], which can converge to the system’s solution faster. Liu et al. proposed a variant of GRK to solve the ridge regression problem [9]. Du Yishu et al. introduced an effective distance-based row selection criterion and proposed the distance-based greedy random Kaczmarz (GDRK) algorithm [10], which has better solving performance than traditional RK and GRK algorithms.
The block Kaczmarz (BK) method [11, 12] is a natural extension of the classical Kaczmarz method and provides another effective way to improve its convergence performance. Unlike the row Kaczmarz method, the block Kaczmarz algorithm involves selecting multiple row indices in each iteration. Given an initial solution $x_{0}$, in the $k$-th iteration, first select a subset of row indices $\tau_{j}\in\left[m\right]$, and then project the iterative solution $x_{j-1}$ onto the solution space $A_{\tau_{j}}x=f_{\tau_{j}}$. Using the pseudo-inverse of $A_{\tau_{j}}$, the Kaczmarz iteration formula for the block is
@@
Recently, Needell et al. proposed the good saving and block random Kaczmarz (RBK) algorithms [13, 14, 15]. Necoara developed a family of randomized block Kaczmarz algorithms that use a subset of constrained and extrapolated steps at each step, and can be deployed on distributed computing units [16]. Subsequent scholars combined RBK with greedy algorithms and proposed the Kaczmarz algorithm for greedy random partitioning (GRBK) [19]. Zhang et al. proposed the greedy random Motzkin Caczmarz (GRMK) method using the greedy selection rule on the maximum residual [17]. Based on the k-means clustering algorithm, Jiang, Li et al. proposed the row clustering block Kaczmarz (RBK(k)) algorithm based on the k-means algorithm [18] and the global random block Kaczmarz method [19], further improving the convergence rate of the algorithm. On this basis, Zheng et al. explored the relationship between ordinary residuals and standardized residuals, and constructed a new random greedy residual block Kaczmarz (GREBK(k)) algorithm [20], which has a much faster computational speed than the RBK(k) method. In 2024, Zhang et al. combined fast projection and weight averaging techniques to propose the aRBK algorithm [21].
In this article, inspired by the GRBK, RBK(k), and GREBK(k) algorithms, the convergence rate of the block iteration Kaczmarz algorithm is indeed much faster than that of the row iteration Kaczmarz algorithm. However, our analysis shows that the sparse structure of the matrix still has a significant impact on the convergence rate of the block Kaczmarz algorithm. For the sparse matrix with a scattered distribution, the iteration time for obtaining an exact solution using the block Kaczmarz algorithm is not ideal. To improve the convergence speed of block Kaczmarz, this paper constructs a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing technology. The proposed algorithm utilizes preprocessing techniques to rearrange the original matrix so that it becomes approximately a banded matrix. Then, the matrix is divided into several blocks, with the goal of grouping orthogonal blocks into one category. In this way, each iteration in this category is performed as two orthogonal projections on nearly orthogonal hyperplanes, so that the iterative solution converges faster to the exact solution. This article theoretically proves the convergence of the POBK algorithm. The numerical experimental results also show that compared with the GRBK, RBK(k), GREBK(k) and aRBK algorithms, the POBK algorithm is more efficient and ideal in terms of algorithm running time.
The structure of this article is as follows: In Section 2, we briefly introduced the GRBK, RBK(k), GREBK(k) and aRBK algorithms. In Section 3, we introduce the preprocessing techniques and partitioning ideas used, analyzed the reasons for their selection, and further proposed a variant POBK algorithm of the block Kaczmarz algorithm. In Section 4, we proved the convergence of the proposed method. In Section 5, we selected some matrices for numerical simulation to demonstrate the effectiveness of our method. Finally, in Section 6, we provide conclusions and point out some future directions for work.
@@ 2 Introduction to existing block methods
The emergence of the block Kaczmarz method has greatly improved the running time and convergence rate of the Kaczmarz method in handling high-dimensional linear systems. The block Kaczmarz method under greedy algorithms and random strategies (GRBK) is a natural development of the block Kaczmarz method. The specific description of the algorithm can be found in Algorithm 1.
Algorithm 1 GRBK Algorithm [18]
1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
3:Randomly obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
4:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
10:end for
In 2022, based on the principle of block Kaczmarz method, Jiang Xianglong et al. first combined k-means clustering algorithm with greedy random technique to propose a novel random block Kaczarz (RBK(k)) method [18]. Based on the distance between the hyperplane formed by each row of the sparse matrix and the current iteration solution, k-means clustering is performed to divide the rows corresponding to the hyperplane with similar distances into one class. In each iteration, the class with the largest distance is selected for orthogonal projection, and the practical probability criterion for selecting the working block submatrix in each iteration is considered an effective variant of the relaxed greedy random Kaczmarz algorithm. The specific description of the RBK(k) algorithm can be found in Algorithm 2.
@@
In 2024, Zheng Wenhao, et al. discovered the relationship between ordinary residuals and standardized residuals [20], that is, the standardized residuals are the 2-norm of the row vector of the corresponding sparse matrix $A$ divided by the ordinary residuals.
$\hat{r}_{i}=\frac{r_{i}}{\left\|A_{(i)}\right\|_{2}}=\frac{f_{i}-A_{(i)}x_{0}}{\left\|A_{(i)}\right\|_{2}}.$ (2.1)
When an initial point is given, the absolute value of the standardized residual is equal to the distance from that initial point to the hyperplanes represented by the linear equation system, i.e.,
$d_{i}=\lvert{\hat{r}_{i}}\rvert$ (2.2)
and the absolute value of standardized residuals is also a measure of greed in the mainstream greedy Kaczmarz algorithm. So they changed the k-means clustering objects in the RBK(k) algorithm from matrix $A$ to standardized residuals, and obtained the partitioning strategy for matrix $A$. They proposed a block Kaczmarz (GREBK(k)) algorithm that combines the k-means algorithm to the block standardized residuals clustering, as shown in Algorithm 3.
@@
3:Apply the k-means method to $d$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
@@
In 2024, Zhang et al. [21] proposed a variant of the random Kaczmarz algorithm, the aRBK algorithm, by combining block projection and weight averaging techniques. Their combination can balance convergence speed, convergence range, and computational complexity. Please refer to Algorithm 4 for more details.
@@
3 Orthogonal Block Kaczmarz Method Based on Preprocessing Technology
3.1 Preprocessing Technology–Sparse Reverse Cuthill-McKee Algorithm
The Kaczmarz algorithm is well known for its good convergence in linear systems, especially for sparse matrices. However, it does not exhibit fast convergence speed for all sparse matrices, particularly those with large dimensions or scattered structural distributions. Even the use of block clustering and greedy algorithms cannot effectively improve this situation. We analyze the iterative characteristics of the Kaczmarz algorithm and the matrix structure in linear systems and find that due to Kaczmarz’s orthogonal projection onto the hyperplane during each iteration, the convergence speed is closely related to the angle between the current iteration point’s hyperplane and the next projection one. It is evident that a smaller angle between these two hyperplanes leads to more iterations and slower convergence, while an angle closer to perpendicular results in fewer iterations and faster convergence.
Taking a two-dimensional linear system as an example, Figure 1 shows the iterative differences between the two cases. $H_{1}$, $H_{2}$, $H_{3}$ represent the three hyperplanes of the linear system, $x_{0}$ represents the initial point, $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ represent the results of Kaczmarz’s first, second, third, and fourth iterations, respectively. The red dots where the hyperplanes intersect represent the exact solution $x_{*}$. The left side of Figure 1 represents the iteration process with a smaller angle between $H_{1}$ and $H_{2}$, and the right side represents the iteration process when $H_{2}$ is perpendicular to $H_{3}$. It can be clearly seen that when two hyperplanes are orthogonal to each other, the Kaczmarz algorithm only needs two steps to reach the exact solution, while the non-orthogonal hyperplanes at this time are still far from the exact solution even after four steps of Kaczmarz iteration. So, we consider whether there is a method to extract rows that are orthogonal to each other in a linear system, or to process the original matrix into a matrix with rows that are orthogonal to each other? If so, the iteration time of the Kaczmarz algorithm will be greatly reduced, see [22].
@@ 3.1 Preprocessing Technology–Sparse Reverse Cuthill-McKee Algorithm
Based on the above analysis, it is considered to be cost-effective to make the original matrix into a matrix whose rows are orthogonal to each other by preprocessing the linear system. Common sparse matrix ordering algorithms include Column Count Reordering, Reverse Cuthill-McKee algorithm (RCM), Minimum Degree Reordering, Nested Dissection and Spectral Reordering, etc. [26, 27, 28, 29, 30, 31]. Each of these algorithms has its own advantages and is suitable for different types of computational tasks. Despite the various methods of ordering available, experimental comparisons have shown that the RCM algorithm usually produces superior computational results in terms of reducing bandwidth and profile of a sparse matrix. The main objective of the RCM algorithm is to minimize the maximum distance between non-zero elements and the main diagonal, thereby reducing bandwidth and greatly improving the orthogonality between matrix row blocks. Consequently, for the block Kaczmarz algorithm, the RCM algorithm is the recommended choice.
The RCM algorithm [26] was initially used for reordering symmetric matrices. It will generate a matrix with rows and columns numbered according to the node numbers. By appropriately renumbering nodes, a new matrix with much smaller matrix bandwidth can usually be generated. The specific principle is that for a symmetric sparse matrix $A$ with a given bandwidth, the RCM algorithm can provide a permutation matrix $P$ such that $\widetilde{A}=PAP^{T}$ has a smaller bandwidth, i.e., the nonzero elements of $\widetilde{A}$ are closer to the diagonal and the run-time complexity is proved to be linear in the number of nonzeros in the matrix [26].
Fortunately, the RCM algorithm also often performs equally well on asymmetric matrices (see http://www.ece.northwestern.edu/local-apps/matlabhelp/techdoc/ref/symrcm.html). Therefore, in our proposed algorithm, we first preprocess the linear system using the RCM algorithm to transform the original coefficient matrix into an approximate banded matrix. Obviously, in the approximate banded matrix, there are many rows or blocks of rows that are orthogonal to each other, see Figure 2, which improves the disadvantage of the random arrangement of the original matrix structure. Based on the above analysis, using the Kaczmarz algorithm again will greatly accelerate the convergence speed.
Figure 2:   The orthogonality representation of the blckhole matrix after being rearranged using the RCM algorithm.
@@
In order to have a clearer and more intuitive view of the processing effect of the RCM algorithm on the structure of sparse matrices, we obtained a series of matrices from the Suite Sparse Matrix Collection (https://sparse.tamu.edu/). Taking matrices poli3, blackhole, dixmaanl, and jagmesh4 as examples, we visualized the sparse structure of these matrices and the matrices that were reordered by the RCM algorithm. We compared the structural changes before and after processing with the RCM algorithm, as shown in Figure 3. The blue part represents the distribution of non-zero elements in the matrix. From Figure 3, it is evident that after being rearranged by the RCM algorithm, the non-zero elements of the matrix are closer to the diagonal. Therefore, choosing the RCM algorithm for preprocessing is reasonable and effective.
@@
Figure 3:   The top four images show the sparse structure of matrices poli3, blackhole, dixmaanl, and jagmesh4, while the bottom four images show the sparse structure of corresponding matrices after RCM rearrangement.
Note that, the run-time complexity of the RCM algorithm is proved to be linear in the number of nonzeros in the matrix [26]. Therefore, the addition of the RCM algorithm does not lead to significant time overhead. We select five matrices muu, poli3, blckhole, kim1, and wathen100 of different dimensions as examples, and count the time consumption of the RCM algorithm for preprocessing them, as shown in Table 1.
$Name$ $m$ $density\left(\%\right)$ $RCM\;preprocessing\;time\;consumption(s)$
$muu$ 7 01270127\,0127 012 $0.337$ 0.002220
$poli3$ 16 9551695516\,95516 955 $0.013$ 0.002261
$blckhole$ 2 13221322\,1322 132 $0.327$ 0.000558
$kim1$ 38 4153841538\,41538 415 $0.063$ 0.018929
$wathen100$ 30 4013040130\,40130 401 $0.051$ 0.005443

### v4

# OBK-RCM: Accelerated Orthogonal Block Kaczmarz Algorithm via RCM Reordering and Dynamic Grouping for Sparse Linear Systems

@@ Abstract
Existing block Kaczmarz methods face challenges in balancing computational efficiency and convergence for large sparse linear systems with scattered nonzero patterns, due to costly partitioning strategies and non-orthogonal projections. In this paper, we propose the orthogonal block Kaczmarz (OBK-RCM) algorithm with the Reverse Cuthill-McKee (RCM), which integrates the RCM reordering with a novel orthogonal block partitioning strategy. RCM transforms sparse matrices into banded structures to enhance inter-block orthogonality, while dynamic grouping of mutually orthogonal blocks based on angle cosine thresholds reduces iterative complexity. In addition, two extended versions (SOBK-RCM and UOBK-RCM) are proposed to deal with non-square systems by constructing extended matrices without sacrificing sparsity. This work offers a practical framework for efficient sparse linear algebra solvers. Experiments on 33 real-world and synthetic matrices show that OBK-RCM achieves 10-50 times faster CPU time (up to several hundred) and 50–90% fewer iterations than state-of-the-art methods (RBK, RBK(k), GREBK(k), aRBK), especially for scattered sparse structures in most cases. Theoretical analysis confirms linear convergence, driven by hyperplane orthogonality.
@@
Kaczmarz algorithm; Sparse linear systems; Reverse Cuthill-McKee ordering; Orthogonal partitioning
@@ 1 Introduction
This paper focuses on solving large-scale sparse linear systems of the form
$Ax=f,$ (1.1)
where $A\in{{\mathbb{R}}^{m\times n}}$, $f\in{{\mathbb{R}}^{m}}$, and $x\in{\mathbb{R}^{n}}$. Nowadays, one of the classic and effective iterative projection methods for solving such systems is the Kaczmarz method [1, 2].
@@
The classic Kaczmarz method suffers from slow convergence due to its dependence on row selection order and lacks rigorous theoretical guarantees. To address these limitations, Strohmer and Vershynin [3, 4] introduced the randomized Kaczmarz (RK) algorithm, achieving an exponential convergence rate in expectation. However, RK’s random row selection criterion remains suboptimal [4]. Subsequent studies proposed greedy variants: Ansorge developed the greedy residual Kaczmarz (GEK) [5], Nutini et al. formulated the greedy distance Kaczmarz (GDK) [6], and Bai and Wu designed the greedy random Kaczmarz (GRK) with relaxation [7, 8]. Further extensions include Liu et al.’s GRK adaptation for ridge regression problem [9] and Du et al.’s distance-based GDRK algorithm [10], which outperforms classical RK and GRK in efficiency.
The block Kaczmarz (BK) method [11, 12] is a natural extension of the classical Kaczmarz method. Unlike the row Kaczmarz method, the block Kaczmarz algorithm involves selecting multiple row indices in each iteration. Given an initial solution $x_{0}$, in the $k$-th iteration, first select a subset of row indices $\tau_{j}\in\left[m\right]$, and then project the iterative solution $x_{j-1}$ onto the solution space $A_{\tau_{j}}x=f_{\tau_{j}}$. Using the pseudo-inverse of $A_{\tau_{j}}$, the Kaczmarz iteration formula for the block is:
@@
Building on these advancements, Needell et al. introduced the randomized block Kaczmarz (RBK) algorithm [13, 14]. Subsequent work integrated greedy strategies into RBK, yielding the greedy random block Kaczmarz (GRBK) method [17]. Zhang et al. further enhanced residual-based selection with the greedy random Motzkin-Kaczmarz (GRMK) algorithm [15], while Jiang and Li incorporated K-means clustering to develop the row-clustered RBK(k) [16] and global randomized block variants [17]. Zheng et al. refined this framework by analyzing residual standardization, proposing the greedy residual block Kaczmarz (GREBK(k)) [18]. Most recently, Zhang et al. (2024) combined fast projection and weighted averaging in the aRBK algorithm [19], achieving a balance between speed and robustness.
While block Kaczmarz methods (e.g., RBK [13], GREBK(k) [18], aRBK [19]) improve convergence over row-wise iterations, two critical limitations hinder their efficiency:
  • 1.
    High partitioning cost: Existing strategies (e.g., K-means clustering [16]) require $O(mk)$ computations for block division, becoming prohibitive for large-scale sparse matrices.
  • 2.
    Slow convergence due to non-orthogonal projections: Random or greedy block selection often results in adjacent blocks with small angles between hyperplanes (Figure 1), leading to redundant iterations.
To address these gaps, we propose the orthogonal block Kaczmarz (OBK-RCM) algorithm with RCM, featuring three core innovations:
  • 1.
    RCM reordering: The Reverse Cuthill-McKee reordering concentrates nonzeros near the diagonal, forming an approximate banded matrix with enhanced inter-block orthogonality (Section 3.1).
  • 2.
    Dynamic orthogonal partitioning: Blocks with near-zero angle cosines are grouped into orthogonal classes (Oclass), enabling rapid convergence via sequential orthogonal projections (Section 3.3).
  • 3.
    Non-square system compatibility: OBK-RCM extends to handle non-square systems without sacrificing sparsity, through extended matrix construction (SOBK-RCM for $m>n$, UOBK-RCM for $m<n$) (see, Eq. 4.1–4.3), avoiding information loss (Section 4.0).
This work bridges the gap between low-cost partitioning and accelerated convergence, offering a unified framework for sparse linear systems.
@@ 2 Introduction to existing block methods
The emergence of the block Kaczmarz method has greatly improved the running time and convergence rate of the Kaczmarz method in handling high-dimensional linear systems.
@@
Firstly, in 2022, Jiang et al. proposed a randomized block Kaczmarz (RBK(k)) method by combining the K-mean clustering algorithm with greedy randomization technique [15], see Algorithm 1 for details.
Secondly, in 2024, based on the relationship between common residuals and standardized residuals, Zheng et al. [17] proposed the block Kaczmarz algorithm (GREBK(k)) for clustering and chunking of standardized residuals combined with the K-means algorithm, see Algorithm 2 for details.
@@
3:Apply the K-means method to the normalized residual vector $d$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
@@
In the same year, Zhang et al. [19] proposed a variant of the random Kaczmarz algorithm, the aRBK algorithm, by combining block projection and weight averaging techniques. Their combination can balance convergence speed, convergence range, and computational complexity. See Algorithm 3 for details.
@@
These algorithms (detailed in Algorithms 1-3) demonstrate how hybrid strategies - clustering of residuals, standardized error metrics, and projection balancing - overcome traditional Kaczmarz limitations in sparse system solving.
3 The orthogonal block Kaczmarz (OBK-RCM) algorithm with RCM
3.1 Matrix reordering to enhance orthogonality
The Kaczmarz algorithm exhibits good convergence properties for linear systems, particularly with sparse matrices. However, its convergence rate slows significantly for large-scale sparse matrices or those with scattered structural distributions, even when enhanced by block clustering and greedy algorithms. This limitation arises because the algorithm’s convergence speed depends critically on the angle between consecutive projection hyperplanes during iterations. Specifically, smaller angles between hyperplanes necessitate more iterations and slower convergence, whereas near-orthogonal projections (angles close to 90°) yield faster convergence with fewer steps. To illustrate, consider a two-dimensional linear system (Figure 1). Here, $H_{1}$, $H_{2}$, $H_{3}$ denote the system’s hyperplanes, $x_{0}$ the initial point, and $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ the iterative solutions. The red dots where the hyperplanes intersect represent the exact solution $x_{*}$. The left panel shows slow convergence when $H_{1}$ and $H_{2}$ form a small angle, while the right panel demonstrates rapid convergence when $H_{1}$ and $H_{3}$ are orthogonal—requiring only two iterations to reach the exact solution $x_{*}$, compared to four for non-orthogonal cases. This motivates the exploration of methods to extract or construct mutually orthogonal rows in the system matrix, which could drastically reduce iteration time (see [22]).
@@ 3.1 Matrix reordering to enhance orthogonality
It is well known that the row blocks of a banded matrix have good orthogonality between them, and the smaller the bandwidth, the stronger the orthogonality. Based on the above analysis, we propose to process the original matrix into an approximate banded matrix by reordering. Sparse matrix reordering algorithms, such as Reverse Cuthill-McKee (RCM), Column Counting (Colperm), Minimum Degree (Amd), Nested Dissection (Dissect)[23, 24, 25, 26], are matrix reduction bandwidth efficient processing techniques. Although there are several sorting methods available, experiments have shown that the RCM algorithm tends to provide better computational results in terms of bandwidth reduction, improved inter-block orthogonality, and algorithmic time-consumption. Therefore, for the block Kaczmarz algorithm, the RCM algorithm is preferred. Section 3.2 describes the RCM algorithm and its advantages in detail.
3.2 Reverse Cuthill-McKee Algorithm
The inverse Cuthill-McKee algorithm [23] is a reordering method for bandwidth minimization of sparse matrices. The algorithm drastically reduces the bandwidth of the matrix and enhances the orthogonality of the blocks by rearranging the order of the columns of the matrix rows and clustering the non-zero elements of the original matrix around the diagonal.
The RCM, Colperm, Amd, and Dissect algorithms have varying results in reducing bandwidth and enhancing orthogonality. We use the matlab built-in function $sprandn$ to randomly generate sparse matrices of different dimensions. Table 1 shows the original bandwidths ($bw$) of these sparse matrices, as well as the bandwidths after reordering by the RCM, Colperm, Amd, and Dissect algorithms. It can be seen that the RCM algorithm can reduce the matrix bandwidth more effectively than the other three sorting algorithms.
Table 1:   Comparison of processing matrix bandwidth of different sorting algorithms.
$Name$ $bw$ $Colperm$ $Amd$ $Dissect$ $RCM$
$1000\times 1000$ 967 984 970 994 28
$5000\times 5000$ 4959 4904 4720 4837 32
$10000\times 10000$ 9779 9813 9769 9514 63
$15000\times 15000$ 14924 14869 14888 14968 68
@@ 3.2 Reverse Cuthill-McKee Algorithm
In order to have a clearer view of the processing effect of the RCM algorithm on the structure of sparse matrices, we obtained a series of matrices from the Suite Sparse Matrix Collection (https://sparse.tamu.edu/). Taking matrices $poli3$, $blckhole$, $dixmaanl$, and $jagmesh4$ as examples, and we visualize the original structure of these matrices and the structure of these matrices after the reordering by the RCM algorithm, see Figure 3, where the blue part is the distribution of non-zero elements of the matrices. It can be clearly seen that after the RCM algorithm reordering, the nonzero elements of the matrices are closer to the diagonal, and the bandwidth of the matrices is significantly reduced.
@@
Figure 2:   The top four images show the sparse structure of matrices poli3, blackhole, dixmaanl, and jagmesh4, while the bottom four images show the sparse structure of corresponding matrices after RCM Reordering.
Meanwhile, the RCM algorithm is simple and efficient with low time complexity. That is, adding the RCM algorithm does not introduce significant time overhead. Taking the matrices $muu$, $poli3$, $rajat07$, $blckhole$, and $jagmesh4$ as an example, the time-consuming statistics of the RCM algorithm for reordering them are shown in the Table 2, where $bw$ denotes the original bandwidth of the matrices, and $bw1$ denotes the bandwidth of the matrices after reordering them by RCM. It can be seen that RCM has obvious advantages in reducing bandwidth and time consumption. For details on the time consumption of the RCM algorithm for reordering sparse matrices, see Ref. [23].
Table 2:    Bandwidth variation and time consumption of the RCM algorithm for reordering sparse matrices.
$Name$ $m$ $density\left(\%\right)$ $bw$ $bw1$ $time~{}consumption(s)$
$muu$ 7102 0.337 4696 311 0.0020
$poli3$ 16955 0.013 16917 5487 0.0022
$rajat07$ 14842 0.029 1483 413 0.0011
$blckhole$ 2132 0.327 1805 105 0.0006
$jagmesh4$ 1440 0.458 1408 56 0.0008
The above results show that the RCM sorting algorithm not only significantly reduces the bandwidth of the matrix, but also has a very low time overhead. Therefore, it is the preferred solution to directly use the RCM sorting algorithm to optimize the structure of sparse linear systems.